Fichas
Probability
Fichas del conjunto (9)
- 1PROBABILITY Number of ways an outcome can happen Total number of possible outcomesIf tossing a coin, what is probability of getting heads?
Number of ways you can get heads = 1
Total possible outcomes = heads and tails = 2
Probability = 1/2 = 50%
When you throw a fair dice there are six possible outcomes: 1, 2, 3, 4, 5 or 6. There are three ways of getting an odd number (1, 3 or 5).
So the probability of getting an odd number equals the number of ways of getting an odd number, divided by the total number of possible outcomes:
Equation: = 3/6 = 1/2 or 50% - 2modemost common average
add all factors together and divide by number of factors - 3mediantake the middle of the factors
- 4modethe most common of the various factors
- 5mutually exclusivemeans that two events cannot both be true at the same time
- 6not mutually exclusivemeans that both events CAN be true at the same time
- 7Example: I have a bag of letters spelling the word MOOSE. If i pick 1 letter out, what is the probability of it being: (a) O (b) a vowel (c) a TANSWER:
(a) 2/5 - very unlikely
(b) 3/5- likely
(c) 0/5- impossible - 8Theoretical possibilitywhen you estimate on the basis of a theory
- 9Experimental probabilitywhen you do an experiment to TEST your theory
Información
- Descripción
- Maths Fichas sobre Probability, creado por ella rose el 18/05/2015.
- Creado por
- ella rose
- Visitas
- 1
- Valoración